Cryan Jeep Tours operates jeep tours in the heart of the Colorado Rockies. The company bases its budgets on two measures of activity (i.e., cost drivers), namely guests and jeeps. One vehicle used in one tour on one day counts as a jeep. Each jeep has one tour guide. The company uses the following data in its budgeting:

Fixed element
per month
Variable element
per guest
Variable element per jeep
Revenue $ 0 $ 277 $ 0
Tour guide wages $ 0 $ 0 $ 155
Vehicle expenses $ 6,200 $ 21 $ 76
Administrative expenses $ 2,700 $ 20 $ 0

In May, the company budgeted for 527 guests and 232 jeeps. The company's income statement showing the actual results for the month appears below:

Cryan Jeep Tours
Income Statement
For the Month Ended May 31
Actual guests 542
Actual jeeps 227
Revenue $ 117,854
Expenses:
Tour guide wages 37,957
Vehicle expenses 23,640
Administrative expenses 10,888
Total expense 72,485
Net operating income $ 45,369

Required:

Prepare a report showing the company's revenue and spending variances for May. Label each variance as favorable (F) or unfavorable (U).
Answer #1
Cryan Jeep Tours
Revenue and spending variances
For the Month Ended May 31
Actual results Flexible Budget Revenue and spending variances
Guests 542 542
Jeeps 227 227
Revenue 117854 150134 32280 U
Expenses:
Tour guide wages 37957 35185 2772 U
Vehicle expenses 23640 34834 11194 F
Administrative expenses 10888 13540 2652 F
Total expense 72485 83559 11074 F
Net operating income 45369 66575 21206 U
Workings:
Flexible Budget
Revenue 150134 =542*277
Expenses:
Tour guide wages 35185 =227*155
Vehicle expenses 34834 =6200+(542*21)+(227*76)
Administrative expenses 13540 =2700+(542*20)